OpenStudy (anonymous):

the curve y=f(x) has a minimum point at (3,5) state the co-ordinates of the corresponding minimum point on the graph of: y=3f(x) and y=f(2x) how do i find the co-ordinates? and what would the f(x) graph look like??

OpenStudy (anonymous):

You don't need to know how the graph looks like in order to get the maximum of y. You can see that the two functions y=3f(x) and y=f(2x) are just f(x) with some operations.

OpenStudy (anonymous):

3f(x) will have same x value as minium , but the y value will be multiplied by 3

OpenStudy (anonymous):

for 3f(x) -> minium ( 3 , 15 ) , methinks , havent taught much about it so might not be right

OpenStudy (anonymous):

For the first function (y=3f(x)), the minimum will occur at the same point x=3, but the value will change by 3 times. That's the minimum point will be (3,15).

OpenStudy (anonymous):

for f(2x) , one thinks use chain rule

OpenStudy (anonymous):

let g(x) = f(2x) g'(x) = 2 f ' (2x) =0 for min/max. f ' (2x) =0 gets a bit strange here

OpenStudy (anonymous):

Well, there is no need to use the chain rule, I guess.

OpenStudy (anonymous):

we know that f' (3) =0 also , but I am not 100% sure if that means 2x =3 , dont think it does

OpenStudy (anonymous):

for the function y=f(2x), the minimum occur at 2x=3 that's x=3/2. The y-value remains the same. The minimum point, then, is (3/2,5).

OpenStudy (anonymous):

but with the second one, shouldn't it stretch the y axis??? and the first one th x axis???

OpenStudy (anonymous):

Thinking of it as "stretching" may lead to confusion, but let's go with that metaphor. - The first one divides y by 3: $$y=3f(x) \implies \frac{y}{3} = f(x) $$ Therefore this is actually a shrinking of the y-axis, relative to the function that says the same size. Since the function stays the same size, everything gets bigger, which is why you multiply by 3 to get the minimum. - The second one multiplies x by 2. This is indeed a stretching of the x-axis, meaning the function gets smaller. That's why you divide the x value by two to get the location of the new minimum.

OpenStudy (anonymous):

ok, thanks :)